OpenStudy (anonymous):

Geometric sequences and series: some guy gets 27,000 $ a year and is expected to receive annual increases of 4%. what will be his salary in his 5th year?

OpenStudy (jdoe0001):

\(\bf {\color{red}{ 4}}\%\ of \ 27,000=\cfrac{{\color{red}{ 4}}}{100}\times 27,000\) what do you think is the common multiplier or factor?

OpenStudy (anonymous):

.04? how would this look in terms of \[\sum_{n}^{x}\]

OpenStudy (jdoe0001):

well, is not a SUM, you're just asked for the 5th year, or 5th term

OpenStudy (anonymous):

right, so what formula would i use?geometric or arithmetic series of a(sub(n)

OpenStudy (jdoe0001):

geometric

OpenStudy (jdoe0001):

hmmm actually... shoot lemme reread

OpenStudy (anonymous):

oops, the starting salary is 32,000$ so: 32000x.04^(5-1)

OpenStudy (anonymous):

that doesnt seem to look right though nor does it give a reasonable answer

OpenStudy (jdoe0001):

right

OpenStudy (anonymous):

i think arithmetic would be correct, 32000+ (5-1)*.04

OpenStudy (anonymous):

nvm

OpenStudy (jdoe0001):

in short, I think this would be the equation for the "continusly compound interest"

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

can't be.... it has to be the compound interest formula so... there =)

OpenStudy (anonymous):

no it has to be sequences, thats what the chapter is on

OpenStudy (jdoe0001):

which means he's getting a 10% increase on his principal compounded yearly thus the compound interest formula \(\bf a_{\color{red}{ t}}=P\left(1+\cfrac{r}{n}\right)^{n{\color{red}{ t}}} \\ \quad \\ n=\textit{periods per year}\qquad t=years\qquad r=rate\qquad P=\textit{initial amount} \\ \quad \\ a_{\color{red}{ 5}}=27000\left(1+\cfrac{0.04}{1}\right)^{1\cdot {\color{red}{ 5}}}\implies a_{\color{red}{ 5}}=27000\left(1+0.04\right)^{{\color{red}{ 5}}}\) the compound interest formulat IS a geometric sequence